Circular No. 5865

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 RO D. Jewitt, University of Hawaii; and J. Luu, Stanford University, report

their discovery of a third faint object with very slow (∼ 3′′/hr) retrograde opposition motion. The object, which appears stellar in 0′′.8 seeing, was detected in CCD images obtained with the University of Hawaii’s 2.2-m telescope on Mauna Kea.

1993 UT  α_{2000}  δ_{2000}  R
Sept. 14.3364  23^{h}22^{m}25.^{s}55  −4^{o}00′52′′.9  23
14.5241  23 22 24.47  −4 00 59.9
15.4216  23 22 19.41  −4 01 33.4
15.4723  23 22 19.12  −4 01 35.1
16.4090  23 22 13.83  −4 02 10.0
16.4216  23 22 13.74  −4 02 10.1
17.4060  23 22 08.18  −4 02 46.8
17.4183  23 22 08.07  −4 02 47.4

Computations by the undersigned indicate that 1993 RO is somewhat closer than the two earlier objects (1992 QB_{1} and 1993 FW). The following is the direct circular orbit that represents the observations, although direct and retrograde orbits of any eccentricity (with shifts by up to ∼ 10 AU in distance) will represent them also. It can be noted that the object is situated ∼ 60^{o} from Neptune.

Epoch = 1993 Sept. 10.00 TT  u = 179.^{o}31
Ω = 170.51 }2000.0
a  = 32.3231 AU  i = 2.53
1993 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Aug. 21 23^{h}24.^{m}66 − 3^{o}46.^{′}2 31.385 32.323 157.^{o}6 0.^{o}7 23.2
31 23 23.77 − 3 52.1 31.337 32.323 167.5 0.4 23.1
Sept. 10 23 22.84 − 3 58.2 31.317 32.323 177.5 0.1 23.1
20 23 21.89 − 4 04.4 31.327 32.323 172.5 0.2 23.1
30 23 20.98 − 4 10.4 31.367 32.323 162.5 0.5 23.1
Oct. 10 23 20.12 − 4 16.0 31.435 32.323 152.4 0.8 23.2
20 23 19.37 − 4 21.0 31.530 32.323 142.3 1.1 23.2
30 23 18.74 − 4 25.1 31.649 32.323 132.1 1.3 23.2
Nov. 9 23 18.26 − 4 28.3 31.788 32.323 122.0 1.5 23.3

1993 September 18 Brian G. Marsden